OpenStudy (anonymous):

compare and contrast the "sum of two squares" and the "difference of two squares

OpenStudy (cwrw238):

difference of 2 squares can be factored sum cannot

OpenStudy (anonymous):

ummmmmmmmmmmmmm............ ok. well, do u know if it would be like x^2+x^2 and x^2-x^2? because if it is, then it would just be adding and subtracting exponents. and yes, it can also be factored, as cwrw238 said.

OpenStudy (cwrw238):

sum can be written in form x^2 + y^2 = (x + y)^2 - 2xy x^2 - y^2 = (x - y)(x + y)

OpenStudy (anonymous):

my question just says Using complete sentences, compare and contrast the "sum of two squares" and the "difference of two squares." Provide an explanation and example of each, including any similarities and differences. thats all i know

OpenStudy (anonymous):

cwrw238 is probably right.